Two similar triangles have perimeters of 15 and 8. If the area of the larger triangle is 100, find the area of the smaller triangle rounded to the nearest tenth.
Found 2 solutions by checkley77, aplused:
Answer by checkley77(12844)

15/100=8/X
15X=800
X=800/15
X=53.3 ANSWER.

Answer by aplused(1)

You must square the ratio of perimeters in order to get the ratio of areas. Therefore:
Ratio of Perimeters: 15/8
Ratio of Areas: 225/64
So, 225/64=100/x
Cross Multiply to get: 6400=225x
Divide by 225 and simplify to get: x=1600/57, which is approximately
27.1 rounded to the nearest tenth